OpenStudy (littlebird):

lim x-> infinity (7x^2 − x + 4)/(3x^2+ 5x − 2)

OpenStudy (danjs):

hello, something different than linear equations for once...

OpenStudy (littlebird):

It'd calculus! ;)

OpenStudy (danjs):

yeah seems about the topic for calc 1 in september

OpenStudy (danjs):

Did you learn the properties of the limit

OpenStudy (littlebird):

sort of. We just started the section, but are allowed to do our homework ahead of time. The quiz for the previous section was earlier today.

OpenStudy (danjs):

As it is, you will not get a value for the limit when you put in infinity for x, you have to try and do a bunch of algebra to change that

OpenStudy (danjs):

the limit properties support this

OpenStudy (littlebird):

I attempted that, but then I got stuck on what algebra to use (factoring failed me)

OpenStudy (danjs):

divide everything by the highest power of x maybe

OpenStudy (littlebird):

ok

OpenStudy (danjs):

recall \[\lim_{x \rightarrow \infty}\frac{ 1 }{ x } = 0\]

OpenStudy (danjs):

as the denominator becomes huge, the fraction 1/huge becomes arbitrarily close to zero

OpenStudy (danjs):

you can see on the graph of y = 1/x for huge x values, it converges to 0

OpenStudy (littlebird):

I got it! :)

OpenStudy (danjs):

7/3

OpenStudy (littlebird):

yes

OpenStudy (danjs):

yeah, that will be on the test prolly.. hah

OpenStudy (danjs):

here is a reference thing

OpenStudy (danjs):

if they really want every notation written, use the properties all the time, like the property 4 there

OpenStudy (littlebird):

thanks

OpenStudy (danjs):

tag me when you get to the Epsilon-Delta Precise definition for the Limit section, to prove a limit exists... I just went back over that recently, fresh in the mind

OpenStudy (danjs):

it is the only section you will prolly not like, or least 90% the class, for some reason

OpenStudy (littlebird):

ok. I hope I'll like it

OpenStudy (danjs):

haha yeah, it is interesting, the only real abstract kind of thing in the course

OpenStudy (danjs):

goodluck, tag me in more if you want,